An insulating sphere is 8.50cm in diameter and carries a 5.85uC charge uniformly distributed throughout its interior volume. Calculate the charge enclosed by a concentric spherical surface with radius r = 2.16cm


1
Expert's answer
2021-01-28T14:01:32-0500

The charge enclosed by a concentric spherical surface with radius r = 2.16 cm can be found by volume it takes:


"v=\frac43\pi r^3."

The volumetric charge density is


"\rho=\frac{Q}{V}=\frac{3Q}{4\pi R^3}."

Then the charge within v will be


"q=\rho v=\frac{3Q}{4\pi R^3}\cdot\frac43\pi r^3=Q\frac{r^3}{R^3}=96\text{ nC}."